OpenStudy (anonymous):

Convert the following equation from polar to rectangular form: theta=pi/3

OpenStudy (amistre64):

..... sqrt(3) that is...

OpenStudy (amistre64):

tan(60) = sqrt(3)/1 y = sqrt(3)x